OpenStudy (anonymous):

Simplify each expression. Use positive exponents to write the answers. 1. (4a^-1b^0)^-3 2. (a^-6/a^-8)^-2 3. (2/3)^-3 4. (3^-2 a^-2 b^12)/(a^4 b^-5)

sam (.sam.):

1. a^(3)/64

sam (.sam.):

2.(a^-6/a^-8)^-2 = (1)/(a^(4)) 3. (2/3)^-3 =27/8 4. (3^-2 a^-2 b^12)/(a^4 b^-5) = b^(17)/9a^(6)

OpenStudy (anonymous):

How?